You drive in a straight line at 24.0 m/s for 10.0 miles, then at 34.0 m/s for another 10.0 miles. Calculate the average speed.
Answer #1

speed = total distance / time 10miles = 16093.4m

total distance  = 10+ 10 miles =20miles = 32186.9

total time =( 16093.4/24) + (16093.4/34)= 670.55+ 473.323 =1143.8735 seconds

speed = 32186.9/ 1143.8735 = 28.138 m/sec